# The numeraire e-variable

### 29 Feb 2024 v1

@@ 3 Duality and reverse information projection
${\mathbb{E}}_{\mathbb{Q}}\left[\log X\right]\leq{\mathbb{E}}_{\mathbb{Q}}\left[-\log\frac{d{\mathbb{P}}^{a}}{d{\mathbb{Q}}}\right]=H({\mathbb{Q}}\mid{\mathbb{P}}).$ (3.7)
Indeed, if $(\log X)^{-}$ has infinite expectation, (3.7) is trivially true. Otherwise we first replace $X$ by $X\wedge n$, in which case (3.6) can be rearranged to (3.7) with $X\wedge n$ in place of $X$. Sending $n$ to infinity and using monotone convergence then gives (3.7). Since equality is achieved by $X=X^{*}$ and ${\mathbb{P}}={\mathbb{P}}^{*}$, we deduce (3.5). ∎
Remark 3.5.
The RIPr ${\mathbb{P}}^{*}$ is not a probability measure in general. Equivalently, it may happen that ${\mathbb{E}}_{\mathbb{Q}}[1/X^{*}]<1$. We will see this phenomenon in Sections 5.1 and 5.4, and other examples are given in Lardy et al. (2023). It is also worth noting that there may exist elements ~$\widetilde{\mathbb{P}}\in{\mathcal{P}}^{\circ\circ}$ that are different from ${\mathbb{P}}^{*}$, but whose absolutely continuous part ~$\widetilde{\mathbb{P}}^{a}$ is equal to ${\mathbb{P}}^{*}$. We will see this in Section 5.4 with $\widetilde{\mathbb{P}}$ being the symmetrization of ${\mathbb{Q}}$. Other examples are obtained from Example 3.6 below.
Example 3.6 (Example 2.2 continued).
Consider a simple null hypothesis $-~{\mathcal{P}}=\{\widetilde{\mathbb{P}}\}$ with $-~{\mathbb{Q}}\ll\widetilde{\mathbb{P}}$ for simplicity so that the numeraire $X^{*}$ is finite. In this case *=~${\mathbb{P}}^{*}=\widetilde{\mathbb{P}}^{a}$ and $-~X^{*}=1/(d{\mathbb{P}}^{*}/d{\mathbb{Q}})=d{\mathbb{Q}}/d\widetilde{\mathbb{P}}$ up to ${\mathbb{Q}}$-nullsets. Here one can easily check that ${\mathcal{P}}^{\circ\circ}$ consists of all sub-probability measures ${\mathbb{P}}$ such that ${\mathbb{P}}(A)\leq\widetilde{\mathbb{P}}(A)$ for all $A$. Indeed, any such ${\mathbb{P}}$ must belong to ${\mathcal{P}}^{\circ\circ}$ since ${\mathbb{E}}_{\mathbb{P}}[X]\leq{\mathbb{E}}_{\widetilde{\mathbb{P}}}[X]\leq 1$ for all e-variables $X$. Conversely, if ${\mathbb{P}}$ is not of this form, then ${\mathbb{P}}(A)>\widetilde{\mathbb{P}}(A)$ for some $A$, and $X=\bm{1}_{A}/\widetilde{\mathbb{P}}(A)$ is an e-variable such that ${\mathbb{E}}_{\mathbb{P}}[X]>1$, showing that ${\mathbb{P}}\notin{\mathcal{P}}^{\circ\circ}$. If ~$\widetilde{\mathbb{P}}(A)=0$ we set $X=\infty\bm{1}_{A}$ in the previous sentence.
We end this section with the general version of Theorem 3.4. The proof of the theorem reduces the general case to the one treated in Theorem 3.4 by introducing the conditional probability measure *${\mathbb{Q}}^{*}(A)={\mathbb{Q}}(A\cap\{X^{*}<\infty\})/{\mathbb{Q}}(X^{*}<\infty)$, while considering the degenerate case ${\mathbb{Q}}(X^{*}<\infty)=0$ separately. We continue to write ${\mathbb{P}}^{a}$ for the absolutely continuous part of ${\mathbb{P}}$ with respect to ${\mathbb{Q}}$, and for a constant $\lambda\in[0,1]$ we write $\lambda\cdot{\mathcal{P}}^{\circ\circ}$ for the set $\{\lambda{\mathbb{P}}\colon{\mathbb{P}}\in{\mathcal{P}}^{\circ\circ}\}\subset{\mathcal{P}}^{\circ\circ}$.
@@
Let $X^{*}$ be the numeraire e-variable and ${\mathbb{P}}^{*}$ be the RIPr, and define the constant $\lambda^{*}={\mathbb{Q}}(X^{*}<\infty)$. Then
[…]
Moreover,
${\mathbb{E}}_{\mathbb{Q}}\left[\frac{d{\mathbb{P}}^{a}}{d{\mathbb{P}}^{*}}\bm{1}_{\{X^{*}<\infty\}}\right]\leq 1\text{ for all }{\mathbb{P}}\in{\mathcal{P}}^{\circ\circ},$ (3.9)
and ${\mathbb{P}}^{*}$ is the unique measure that satisfies (3.8) and (3.9). Furthermore, (3.9) is equivalent to
${\mathbb{E}}_{\mathbb{Q}}\left[\log\left(\lambda^{*}\frac{d{\mathbb{P}}^{a}}{d{\mathbb{P}}^{*}}\right)\bm{1}_{\{X^{*}<\infty\}}\right]\leq 0\text{ for all }{\mathbb{P}}\in{\mathcal{P}}^{\circ\circ}.$ (3.10)
Finally, one still has the strong duality relation (3.5).
@@
We first consider the fully degenerate case $\lambda^{*}={\mathbb{Q}}(X^{*}<\infty)=0$ and hence *${\mathbb{P}}^{*}=0$. In this case (3.9) and (3.10) reduce to the statements that $0\leq 1$ and $0\leq 0$, which are true and hence trivially equivalent. The uniqueness statement follows because $\lambda^{*}\cdot{\mathcal{P}}^{\circ\circ}=\{0\}$, so only the zero measure can satisfy (3.8). The strong duality (3.5) holds because all four terms are infinite.
Next, in the case $\lambda^{*}=1$, (3.8) states that ${\mathbb{P}}^{*}$ is equivalent to ${\mathbb{Q}}$ and belongs to ${\mathcal{P}}^{\circ\circ}$, which we already know from Lemma 3.1. Thus the theorem simply reduces to Theorem 3.4 in this case.
Finally, suppose $\lambda^{*}\in(0,1)$. We first note that the strong duality (3.5) holds because all four terms are infinite. Next, we verify (3.8). The stated equivalence on $\{X^{*}<\infty\}$ is evident from the definition of ${\mathbb{P}}^{*}$. To show that ${\mathbb{P}}^{*}$ belongs to $\lambda^{*}\cdot{\mathcal{P}}^{\circ\circ}$ we refine the calculation in the proof of Lemma 3.1. Consider any $X\in{\mathcal{E}}$ and set $X^{\prime}=X\bm{1}_{\{X^{*}<\infty\}}+\infty\bm{1}_{\{X^{*}=\infty\}}$. This still belongs to ${\mathcal{E}}$ since it ${\mathbb{P}}$-almost surely equals $X$ for all ${\mathbb{P}}\in{\mathcal{P}}$. Thanks to the convention (1.1) we have $X^{\prime}/X^{*}=\bm{1}_{\{X^{*}<\infty\}}X^{\prime}/X^{*}+\bm{1}_{\{X^{*}=\infty\}}$ and hence
${\mathbb{E}}_{{\mathbb{P}}^{*}}[X]={\mathbb{E}}_{{\mathbb{P}}^{*}}[X^{\prime}]={\mathbb{E}}_{\mathbb{Q}}\left[\bm{1}_{\{X^{*}<\infty\}}\frac{X^{\prime}}{X^{*}}\right]={\mathbb{E}}_{\mathbb{Q}}\left[\frac{X^{\prime}}{X^{*}}\right]-{\mathbb{Q}}(X^{*}=\infty)\leq{\mathbb{Q}}(X^{*}<\infty),$
using the numeraire property in the last step. It follows that *${\mathbb{P}}^{*}\in\lambda^{*}\cdot{\mathcal{P}}^{\circ\circ}$.
We now introduce the conditional probability measures *${\mathbb{Q}}^{*}(A)={\mathbb{Q}}(A\cap\{X^{*}<\infty\})/\lambda^{*}$ and **${\mathbb{P}}^{**}(A)={\mathbb{P}}^{*}(A)/\lambda^{*}$. Taking ${\mathbb{Q}}^{*}$ as our new alternative hypothesis, we have that $X^{*}$ is still the numeraire. Indeed, it is a ${\mathbb{Q}}^{*}$-almost surely strictly positive e-variable, and for any other e-variable $X$ we define $X^{\prime}$ as above and get
${\mathbb{E}}_{{\mathbb{Q}}^{*}}\left[\frac{X}{X^{*}}\right]=\frac{1}{\lambda^{*}}{\mathbb{E}}_{\mathbb{Q}}\left[\bm{1}_{\{X^{*}<\infty\}}\frac{X}{X^{*}}\right]=\frac{1}{\lambda^{*}}\left({\mathbb{E}}_{\mathbb{Q}}\left[\frac{X^{\prime}}{X^{*}}\right]-{\mathbb{Q}}(X^{*}=\infty)\right)\leq 1.$
Since *${\mathbb{Q}}^{*}(X^{*}<\infty)=1$ by construction, we have from Theorem 2.5 that *${\mathbb{Q}}^{*}\ll{\mathcal{P}}$. We also have $d{\mathbb{P}}^{**}/d{\mathbb{Q}}^{*}=1/X^{*}$, so ${\mathbb{P}}^{**}$ is the RIPr corresponding to the alternative hypothesis ${\mathbb{Q}}^{*}$. Now, observe that (3.9) and (3.10) can be expressed in terms of ${\mathbb{Q}}^{*}$ and ${\mathbb{P}}^{**}$, and become
${\mathbb{E}}_{{\mathbb{Q}}^{*}}\left[\frac{d{\mathbb{P}}^{a}}{d{\mathbb{P}}^{**}}\right]\leq 1\text{ for all }{\mathbb{P}}\in{\mathcal{P}}^{\circ\circ}$
and
${\mathbb{E}}_{{\mathbb{Q}}^{*}}\left[\log\frac{d{\mathbb{P}}^{a}}{d{\mathbb{P}}^{**}}\right]\leq 0\text{ for all }{\mathbb{P}}\in{\mathcal{P}}^{\circ\circ}.$
We are now in a position to apply Theorem 3.4 with ${\mathbb{Q}}^{*}$ and ${\mathbb{P}}^{**}$ in place of ${\mathbb{Q}}$ and ${\mathbb{P}}^{*}$. We deduce that (3.9) and (3.10) hold, that they are equivalent whenever ${\mathbb{P}}^{**}$ is an element of ${\mathcal{P}}^{\circ\circ}$ equivalent to ${\mathbb{Q}}^{*}$, and that ${\mathbb{P}}^{**}$ is the unique such element that satisfies (3.9). Since ${\mathbb{P}}^{**}$ is equivalent to ${\mathbb{Q}}^{*}$ and belongs to ${\mathcal{P}}^{\circ\circ}$ if and only if (3.8) holds, the proof of the theorem is complete. ∎
4 Finding the numeraire
So far we have studied the abstract existence, uniqueness, and duality theory of the numeraire, but we have not yet said much about how to actually compute it. In this section we develop tools for doing so, which we then apply to several examples in Section 5. We focus on the statistically most relevant situation where ${\mathbb{Q}}\ll{\mathcal{P}}$. In this case the numeraire $X^{*}$, and indeed every e-variable, is ${\mathbb{Q}}$-almost surely finite and the RIPr ${\mathbb{P}}^{*}$ is equivalent to ${\mathbb{Q}}$.
Proposition 2.4 suggests that $X^{*}$ can be found by solving an optimization problem. Unfortunately this may not be straightforward because the optimization is over the set of all e-variables, which can be difficult to characterize. For the same reason, even if a candidate $X^{*}$ has been found it may not be clear how to check that it does, in fact, satisfy the numeraire property. Alternatively, the strong duality of Theorem 3.4 suggests minimizing relative entropy to find ${\mathbb{P}}^{*}$. However, the optimization is again over a set that can be difficult to describe explicitly, in this case the effective null ${\mathcal{P}}^{\circ\circ}$. Moreover, checking that a candidate ${\mathbb{P}}^{*}$ is optimal involves comparing it to all elements of ${\mathcal{P}}^{\circ\circ}$.
We first present a verification theorem which simplifies the task of checking that candidates $X^{*}$ and ${\mathbb{P}}^{*}$ are in fact optimal. Despite the almost trivial proof, the result turns out to be remarkably useful.
Theorem 4.1.
Assume that ${\mathbb{Q}}\ll{\mathcal{P}}$. Let $X^{*}$ be an e-variable that is ${\mathbb{Q}}$-almost surely strictly positive, and let ${\mathbb{P}}^{*}$ be given by $d{\mathbb{P}}^{*}/d{\mathbb{Q}}=1/X^{*}$. Then $X^{*}$ is a numeraire if and only if ${\mathbb{P}}^{*}$ belongs to the effective null ${\mathcal{P}}^{\circ\circ}$. Thus, the numeraire is the only e-variable that is also a likelihood ratio between ${\mathbb{Q}}$ and some equivalent element of ${\mathcal{P}}^{\circ\circ}$.
Proof.
@@ 4 Finding the numeraire
If $X^{*}$ is a numeraire, then Lemma 3.1 yields that *${\mathbb{P}}^{*}\in{\mathcal{P}}^{\circ\circ}$. Let us now argue the converse. The definition of ${\mathbb{P}}^{*}$ and the fact that it belongs to ${\mathcal{P}}^{\circ\circ}$ yield ${\mathbb{E}}_{\mathbb{Q}}[X/X^{*}]={\mathbb{E}}_{{\mathbb{P}}^{*}}[X]\leq 1$ for any e-variable $X$. This is the numeraire property. ∎
@@
The following corollary applies in settings where the null hypothesis is ‘generated’, in the sense of (4.1) below, by a subset ${\mathcal{E}}_{0}$ of e-variables. It shows that once a candidate $X^{*}$ has been found, it is enough to check the numeraire property for the generating family ${\mathcal{E}}_{0}$, provided that it holds with equality for the e-variable which is identically one. The latter condition means that the sub-probability measure ${\mathbb{P}}^{*}$ is in fact a probability measure.
@@
Assume that ${\mathbb{Q}}\ll{\mathcal{P}}$. Let ${\mathcal{E}}_{0}$ be a family of $[0,\infty]$-valued random variables that generates the null hypothesis in the sense that
${\mathcal{P}}=\{{\mathbb{P}}\in M_{1}\colon{\mathbb{E}}_{\mathbb{P}}[X]\leq 1\text{ for all }X\in{\mathcal{E}}_{0}\}.$ (4.1)
Let $X^{*}$ be a ${\mathbb{Q}}$-almost surely strictly positive e-variable such that
${\mathbb{E}}_{\mathbb{Q}}\left[\frac{1}{X^{*}}\right]=1\text{ and }{\mathbb{E}}_{\mathbb{Q}}\left[\frac{X}{X^{*}}\right]\leq 1\text{ for all }X\in{\mathcal{E}}_{0}.$ (4.2)
@@
Let ${\mathbb{P}}^{*}$ be given by $d{\mathbb{P}}^{*}/d{\mathbb{Q}}=1/X^{*}$. Then (4.2) says that ${\mathbb{P}}^{*}$ is a probability measure such that ${\mathbb{E}}_{{\mathbb{P}}^{*}}[X]\leq 1$ for all $X\in{\mathcal{E}}_{0}$. Thus by (4.1), ${\mathbb{P}}^{*}$ belongs to ${\mathcal{P}}$, hence to ${\mathcal{P}}^{\circ\circ}$. The result now follows from Theorem 4.1. ∎
@@
Consider the single coin toss space $\Omega=\{0,1\}$ and the random variable $X_{0}$ given by $X_{0}(0)=2$ and $X_{0}(1)=1/2$. Let the singleton set ${\mathcal{E}}_{0}=\{X_{0}\}$ generate the null hypothesis ${\mathcal{P}}$, which then consists of all probability measures ${\mathbb{P}}$ with ${\mathbb{P}}(0)\leq 1/3$. Next, let the alternative ${\mathbb{Q}}$ be given by ${\mathbb{Q}}(0)=1$, ${\mathbb{Q}}(1)=0$. We then have ${\mathbb{E}}_{\mathbb{Q}}[1/X_{0}]=1/2<1$ and, trivially, ${\mathbb{E}}_{\mathbb{Q}}[X/X_{0}]\leq 1$ for $X\in{\mathcal{E}}_{0}$. If Corollary 4.2 were to remain true with an inequality instead of an equality in (4.2), we would conclude that $X_{0}$ is the numeraire. However, this is not so. Indeed, for the e-variable $X^{*}$ given by $X^{*}(0)=3$, $X^{*}(1)=0$, we have ${\mathbb{E}}_{\mathbb{Q}}[X^{*}/X_{0}]=3/2>1$. As suggested by the notation, $X^{*}$ is in fact the numeraire in this example. To see this, just note that every e-variable $X$ must satisfy $X(0)\leq 3$ since ${\mathcal{P}}$ contains the probability measure ${\mathbb{P}}$ with ${\mathbb{P}}(0)=1/3$, and hence ${\mathbb{E}}_{\mathbb{Q}}[X/X^{*}]=X(0)/3\leq 1$. It is also worth noting that ${\mathbb{E}}_{\mathbb{Q}}[1/X^{*}]=1/3<1$, which means that we cannot use Corollary 4.2 to detect the numeraire property of $X^{*}$. The corollary is not a universal tool, but nonetheless a useful one as we will see in Section 5.
Next, we consider the dual approach of finding ${\mathbb{P}}^{*}$ directly, either by minimizing relative entropy or, if a candidate is available, checking that (3.3) or (3.4) is satisfied. The difficulty is that the domain is all of ${\mathcal{P}}^{\circ\circ}$, which can be hard to describe explicitly. Matters would become simpler if one could instead work with ${\mathcal{P}}$. We first show that this is not possible in general.
@@
This example shows that (3.3) may no longer characterize ${\mathbb{P}}^{*}$ uniquely if ${\mathcal{P}}^{\circ\circ}$ is replaced by ${\mathcal{P}}$. Let $\Omega=[0,1]$ with its Borel $\sigma$-algebra, and let ${\mathcal{P}}=\{\delta_{\omega}\colon\omega\in[0,1]\}$ be the set of all Dirac point masses. In this case the set of e-variables is trivial in the sense that a random variable $X$ is an e-variable if and only if $0\leq X(\omega)\leq 1$ for all $\omega\in[0,1]$. It follows that ${\mathcal{P}}^{\circ\circ}$ contains all sub-probabilities on $\Omega$. Now, let the alternative hypothesis ${\mathbb{Q}}$ be the uniform distribution on the unit interval. We then have the extreme situation that ${\mathcal{P}}$ does not contain any distribution absolutely continuous with respect to ${\mathbb{Q}}$, whereas ${\mathcal{P}}^{\circ\circ}$ contains all such distributions. In particular, (3.3) would be vacuously true for any candidate ${\mathbb{P}}^{*}$ if ${\mathcal{P}}^{\circ\circ}$ were replaced by ${\mathcal{P}}$. In contrast, as currently stated, (3.3) will only hold for ${\mathbb{P}}^{*}={\mathbb{Q}}$, which is the correct answer.
The situation improves in the presence of a reference measure $\mu$ with respect to which ${\mathbb{Q}}$ and all ${\mathbb{P}}\in{\mathcal{P}}$ are absolutely continuous. In this setting, if ${\mathbb{P}}_{n},{\mathbb{P}}$ are measures with densities $p_{n}=d{\mathbb{P}}_{n}/d\mu$ and $p=d{\mathbb{P}}/d\mu$ we say that ${\mathbb{P}}_{n}$ $\mu$-converges to ${\mathbb{P}}$ if $p_{n}$ converges to $p$ in probability under $\mu$. A set $C$ of absolutely continuous measures is called $\mu$-closed if it is closed with respect to $\mu$-convergence. It is called solid if it contains every (nonnegative) measure ${\mathbb{P}}^{\prime}$ that is dominated by some ${\mathbb{P}}\in C$ in the sense that ${\mathbb{P}}^{\prime}(A)\leq{\mathbb{P}}(A)$ for all $A$. The following result characterizes ${\mathcal{P}}^{\circ\circ}$ in terms of these concepts.
@@
Suppose all ${\mathbb{P}}\in{\mathcal{P}}$ are absolutely continuous with respect to a probability measure $\mu$. Then all elements of ${\mathcal{P}}^{\circ\circ}$ are also absolutely continuous with respect to $\mu$, and ${\mathcal{P}}^{\circ\circ}$ is the smallest $\mu$-closed convex solid set that contains ${\mathcal{P}}$.

### 22 Mar 2024 v2

@@ 3 Duality and reverse information projection
${\mathbb{E}}_{{\mathsf{Q}}^{*}}\left[\frac{d{\mathsf{P}}^{a}}{d{\mathsf{P}}^{**}}\right]\leq 1\text{ for all }{\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}\quad\text{and}\quad{\mathbb{E}}_{{\mathsf{Q}}^{*}}\left[\log\frac{d{\mathsf{P}}^{a}}{d{\mathsf{P}}^{**}}\right]\leq 0\text{ for all }{\mathsf{P}}\in{\mathcal{P}}^{\circ\circ},$
respectively. We may thus simply apply Theorem 3.4 with ${\mathsf{Q}}^{*}$ and ${\mathsf{P}}^{**}$ in place of ${\mathsf{Q}}$ and ${\mathsf{P}}^{*}$ to deduce that the three statements are indeed equivalent and that the ‘conditional’ strong duality (3.7) holds. The ‘unconditional’ strong duality (3.3) holds because all four terms are infinite. ∎
Following up on Example 3.7, the following corollary records the RIPr and numeraire for a point null, generalizing a well known result that the likelihood ratio is the optimal e-variable if the two distributions are equivalent; see for example Shafer (2021).
Corollary 3.9.
Consider an arbitrary point null ${\mathcal{P}}=\{{\mathsf{P}}_{0}\}$. The RIPr is the absolutely continuous part ${\mathsf{P}}^{*}={\mathsf{P}}_{0}^{a}$ with respect to ${\mathsf{Q}}$, and the numeraire is
$X^{*}=\infty\bm{1}_{N}+\frac{d{\mathsf{Q}}^{*}}{d{\mathsf{P}}_{0}}\bm{1}_{N^{c}},$ (3.8)
where $N$ is a set such that ${\mathsf{P}}_{0}(N)=0$ and ${\mathsf{Q}}$ restricted to $N^{c}$ is absolutely continuous with respect to ${\mathsf{P}}_{0}$, and ${\mathsf{Q}}^{*}={\mathsf{Q}}(\cdot\mid N^{c})$ whenever ${\mathsf{Q}}(N^{c})>0$.
Proof.
If ${\mathsf{P}}_{0}$ and ${\mathsf{Q}}$ are singular we have ${\mathsf{P}}^{*}=0$ and $X^{*}=\infty$ which are indeed the RIPr and numeraire, so we focus on the non-singular case. We first compute the numeraire. Thanks to Example 3.7, the numeraire for ${\mathsf{Q}}^{*}$ is $d{\mathsf{Q}}^{*}/d{\mathsf{P}}_{0}$ up to ${\mathsf{Q}}^{*}$-nullsets. The numeraire for ${\mathsf{Q}}$ should be infinite on $N$, which leads to (3.8). We note that $X^{*}$ is ${\mathsf{Q}}$-almost surely strictly positive, it is an e-variable because ${\mathbb{E}}_{{\mathsf{P}}_{0}}[X^{*}]=1$, and for any other e-variable $X$ we have ${\mathbb{E}}_{\mathsf{Q}}[X/X^{*}]\leq{\mathsf{Q}}(N)+{\mathsf{Q}}(N^{c}){\mathbb{E}}_{{\mathsf{Q}}^{*}}[X/(d{\mathsf{Q}}^{*}/d{\mathsf{P}}_{0})]\leq 1$. This confirms that $X^{*}$ is the numeraire, and since $1/X^{*}=d{\mathsf{P}}_{0}^{a}/d{\mathsf{Q}}$ up to ${\mathsf{Q}}$-nullsets the RIPr is ${\mathsf{P}}^{*}={\mathsf{P}}_{0}^{a}$ as claimed. ∎
4 Finding the numeraire
So far we have studied the abstract existence, uniqueness, and duality theory of the numeraire, but we have not yet said much about how to actually compute it. In this section we develop tools for doing so, which we then apply to several examples in Section 5. We focus on the statistically most relevant situation where ${\mathsf{Q}}\ll{\mathcal{P}}$. In this case the numeraire $X^{*}$, and indeed every e-variable, is ${\mathsf{Q}}$-almost surely finite and the RIPr ${\mathsf{P}}^{*}$ is equivalent to ${\mathsf{Q}}$. The general case can be reduced to this case by replacing ${\mathsf{Q}}$ with the conditional measure ${\mathsf{Q}}^{*}$ in Theorem 3.8 (except in the fully degenerate situation where the numeraire is infinity and the RIPr zero.)
Proposition 2.4 suggests that $X^{*}$ can be found by solving an optimization problem. Unfortunately this may not be straightforward because the optimization is over the set of all e-variables, which can be difficult to characterize. For the same reason, even if a candidate $X^{*}$ has been found it may not be clear how to check that it does, in fact, satisfy the numeraire property. Alternatively, the strong duality of Theorem 3.4 suggests minimizing relative entropy to find ${\mathsf{P}}^{*}$. However, the optimization is again over a set that can be difficult to describe explicitly, in this case the effective null ${\mathcal{P}}^{\circ\circ}$. Moreover, checking that a candidate ${\mathsf{P}}^{*}$ is optimal involves comparing it to all elements of ${\mathcal{P}}^{\circ\circ}$.
We first present a verification theorem which simplifies the task of checking that candidates $X^{*}$ and ${\mathsf{P}}^{*}$ are in fact optimal. Despite the simple proof, the result turns out to be remarkably useful.
@@ 4 Finding the numeraire
Assume that ${\mathsf{Q}}\ll{\mathcal{P}}$. Let $X^{*}$ be an e-variable that is ${\mathsf{Q}}$-almost surely strictly positive and let ${\mathsf{P}}^{*}$ be given by $d{\mathsf{P}}^{*}/d{\mathsf{Q}}=1/X^{*}$. Then $X^{*}$ is a numeraire if and only if ${\mathsf{P}}^{*}$ belongs to the effective null ${\mathcal{P}}^{\circ\circ}$. Thus the numeraire is the only e-variable that is also a likelihood ratio between ${\mathsf{Q}}$ and some equivalent element of ${\mathcal{P}}^{\circ\circ}$. Finally, ${\mathbb{E}}_{{\mathsf{P}}^{*}}[X^{*}]=1$ and ${\mathsf{P}}^{*}$ is maximal in the sense that no other ${\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}$ equivalent to ${\mathsf{Q}}$ can satisfy ${\mathsf{P}}(A)\geq{\mathsf{P}}^{*}(A)$ for all $A$ with strict inequality for some $A$.
@@
Recall that all e-variables are ${\mathsf{Q}}$-almost surely finite thanks to Theorem 2.5 and the assumption that ${\mathsf{Q}}\ll{\mathcal{P}}$; in particular, ${\mathsf{P}}^{*}$ is equivalent to ${\mathsf{Q}}$. Now, if $X^{*}$ is a numeraire, then Lemma 3.1 yields that ${\mathsf{P}}^{*}\in{\mathcal{P}}^{\circ\circ}$. For the converse, the definition of ${\mathsf{P}}^{*}$ and the fact that it belongs to ${\mathcal{P}}^{\circ\circ}$ yield ${\mathbb{E}}_{\mathsf{Q}}[X/X^{*}]={\mathbb{E}}_{{\mathsf{P}}^{*}}[X]\leq 1$ for any e-variable $X$. This is the numeraire property. Finally, the definition of ${\mathsf{P}}^{*}$ immediately gives ${\mathbb{E}}_{{\mathsf{P}}^{*}}[X^{*}]=1$, and ${\mathsf{P}}^{*}$ must be maximal because if a ‘dominating’ equivalent ${\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}$ existed we would get the contradiction $1={\mathbb{E}}_{{\mathsf{P}}^{*}}[X^{*}]<{\mathbb{E}}_{\mathsf{P}}[X^{*}]\leq 1$. ∎
@@
The following corollary applies in settings where the null hypothesis is ‘generated’, in the sense of (4.1) below, by a subset ${\mathcal{E}}_{0}$ of e-variables. Null hypotheses of this kind were studied by Casgrain et al. (2024) in a sequential setting. The corollary shows that once a candidate $X^{*}$ has been found, it is enough to check the numeraire property for the generating family ${\mathcal{E}}_{0}$, provided that it holds with equality for the e-variable which is identically one. The latter condition means that the RIPr ${\mathsf{P}}^{*}$ is a probability measure.
@@
Assume that ${\mathsf{Q}}\ll{\mathcal{P}}$. Let ${\mathcal{E}}_{0}$ be a family of $[0,\infty]$-valued random variables that generates the null hypothesis in the sense that
${\mathcal{P}}=\{{\mathsf{P}}\in M_{1}\colon{\mathbb{E}}_{\mathsf{P}}[X]\leq 1\text{ for all }X\in{\mathcal{E}}_{0}\}.$ (4.1)
Let $X^{*}$ be a ${\mathsf{Q}}$-almost surely strictly positive e-variable such that
${\mathbb{E}}_{\mathsf{Q}}\left[\frac{1}{X^{*}}\right]=1\text{ and }{\mathbb{E}}_{\mathsf{Q}}\left[\frac{X}{X^{*}}\right]\leq 1\text{ for all }X\in{\mathcal{E}}_{0}.$ (4.2)
@@
Let ${\mathsf{P}}^{*}$ be given by $d{\mathsf{P}}^{*}/d{\mathsf{Q}}=1/X^{*}$. Then (4.2) says that ${\mathsf{P}}^{*}$ is a probability measure such that ${\mathbb{E}}_{{\mathsf{P}}^{*}}[X]\leq 1$ for all $X\in{\mathcal{E}}_{0}$. Thus by (4.1), ${\mathsf{P}}^{*}$ belongs to ${\mathcal{P}}$, hence to ${\mathcal{P}}^{\circ\circ}$. The result now follows from Theorem 4.1. ∎
@@
Consider the single coin toss space $\Omega=\{0,1\}$ and the random variable $X_{0}$ given by $X_{0}(0)=2$ and $X_{0}(1)=1/2$. Let the singleton set ${\mathcal{E}}_{0}=\{X_{0}\}$ generate the null hypothesis ${\mathcal{P}}$, which then consists of all probability measures ${\mathsf{P}}$ with ${\mathsf{P}}(0)\leq 1/3$. Next, let the alternative ${\mathsf{Q}}$ be given by ${\mathsf{Q}}(0)=1$, ${\mathsf{Q}}(1)=0$. We then have ${\mathbb{E}}_{\mathsf{Q}}[1/X_{0}]=1/2<1$ and, trivially, ${\mathbb{E}}_{\mathsf{Q}}[X/X_{0}]\leq 1$ for $X\in{\mathcal{E}}_{0}$. If Corollary 4.2 were to remain true with an inequality instead of an equality in (4.2), we would conclude that $X_{0}$ is the numeraire. However, this is not so. Indeed, for the e-variable $X^{*}$ given by $X^{*}(0)=3$, $X^{*}(1)=0$, we have ${\mathbb{E}}_{\mathsf{Q}}[X^{*}/X_{0}]=3/2>1$. As suggested by the notation, $X^{*}$ is in fact the numeraire in this example. To see this, just note that every e-variable $X$ must satisfy $X(0)\leq 3$ since ${\mathcal{P}}$ contains the probability measure ${\mathsf{P}}$ with ${\mathsf{P}}(0)=1/3$, and hence ${\mathbb{E}}_{\mathsf{Q}}[X/X^{*}]=X(0)/3\leq 1$. It is also worth noting that ${\mathbb{E}}_{\mathsf{Q}}[1/X^{*}]=1/3<1$, which means that we cannot use Corollary 4.2 to detect the numeraire property of $X^{*}$. The corollary is not a universal tool, but nonetheless a useful one as we will see in Section 5.
Next, we consider the dual approach of finding ${\mathsf{P}}^{*}$ directly, either by minimizing relative entropy or, if a candidate is available, checking that (ii) or (iii) of Theorem 3.4 are satisfied. The difficulty is that the domain is all of ${\mathcal{P}}^{\circ\circ}$, which can be hard to describe explicitly. Matters would become simpler if one could instead work with ${\mathcal{P}}$. We first show that this is not possible in general.
@@
This example shows that Theorem 3.4(ii) may no longer characterize the RIPr ${\mathsf{P}}^{*}$ uniquely if ${\mathcal{P}}^{\circ\circ}$ is replaced by ${\mathcal{P}}$. Let $\Omega=[0,1]$ with its Borel $\sigma$-algebra, and let ${\mathcal{P}}=\{\delta_{\omega}\colon\omega\in[0,1]\}$ be the set of all Dirac point masses. In this case the set of e-variables is trivial in the sense that a random variable $X$ is an e-variable if and only if $0\leq X(\omega)\leq 1$ for all $\omega\in[0,1]$. It follows that ${\mathcal{P}}^{\circ\circ}$ contains all sub-probabilities on $\Omega$. Now, let the alternative hypothesis ${\mathsf{Q}}$ be the uniform distribution on the unit interval. We then have the extreme situation that ${\mathcal{P}}$ does not contain any distribution absolutely continuous with respect to ${\mathsf{Q}}$, whereas ${\mathcal{P}}^{\circ\circ}$ contains all such distributions. In particular, Theorem 3.4(ii) would be vacuously true for any candidate ${\mathsf{P}}^{*}$ if ${\mathcal{P}}^{\circ\circ}$ were replaced by ${\mathcal{P}}$. In contrast, as currently stated, Theorem 3.4(ii) will only hold for ${\mathsf{P}}^{*}={\mathsf{Q}}$, which is the correct answer.
The situation improves in the presence of a reference measure $\mu$ with respect to which ${\mathsf{Q}}$ and all ${\mathsf{P}}\in{\mathcal{P}}$ are absolutely continuous. In this setting, if ${\mathsf{P}}_{n},{\mathsf{P}}$ are measures with densities $p_{n}=d{\mathsf{P}}_{n}/d\mu$ and $p=d{\mathsf{P}}/d\mu$ we say that ${\mathsf{P}}_{n}$ $\mu$-converges to ${\mathsf{P}}$ if $p_{n}$ converges to $p$ in probability under $\mu$. A set $C$ of absolutely continuous measures is called $\mu$-closed if it is closed with respect to $\mu$-convergence. It is called solid if it contains every (nonnegative) measure ${\mathsf{P}}^{\prime}$ that is dominated by some ${\mathsf{P}}\in C$ in the sense that ${\mathsf{P}}^{\prime}(A)\leq{\mathsf{P}}(A)$ for all $A$. The following result characterizes ${\mathcal{P}}^{\circ\circ}$ in terms of these concepts.
@@
Suppose all ${\mathsf{P}}\in{\mathcal{P}}$ are absolutely continuous with respect to a probability measure $\mu$. Then all elements of ${\mathcal{P}}^{\circ\circ}$ are also absolutely continuous with respect to $\mu$, and ${\mathcal{P}}^{\circ\circ}$ is the smallest $\mu$-closed convex solid set that contains ${\mathcal{P}}$.